OpenStudy (anonymous):

calculus implicit differentiation x^2 + y + 2y^3 = 0

OpenStudy (anonymous):

what do you want to find dy/dx?

OpenStudy (anonymous):

yess

OpenStudy (lalaly):

\[2x+ \frac{dy}{dx}+6y^2 \frac{dy}{dx}=0\]

OpenStudy (anonymous):

dy/dx=-(2x)/(1+6y^2)

OpenStudy (anonymous):

\[x^{2} + y + 2y^{3} = 0\]\[2x(dx) + dy + 6y^{2}(dy) = 0\]\[\frac{dy}{dx} = \frac{-2x}{1 + 6y^{2}}\]